# The physical basis of teleological evolutions

### v12

@@ 1 Rationale
This work proposes a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the very goal it will produce in the future. It is, of course, a puzzling notion since it implies a causal loop. For more than two millennia, it was applied by natural philosophy first to the functional structure of the living (Aristotle: teeth were created for the purpose of chewing) and then, with the advent of the notion of evolution of the living, to these evolutions. It was then dismissed by modern science because of its alleged lack of physical basis.
What we believe should change this (centuries-old) position of science is the advent of quantum mechanics and, more specifically, the relatively recent discovery of the quantum computational speedup. In fact, we will show that quantum algorithms are scientific examples of teleological evolutions: it is their teleological character that yields their quantum computational speedup over the corresponding classical algorithms.
This is not just the rehabilitation by science of an abandoned philosophical notion. Once represented in a scientific way, the teleological notion fires back at mainstream science by challenging its current rejection of the notion of retrocausality.
To explain in sufficient detail from the beginning the present unconventional approach to quantum computation, to the usual “Introduction” we have preferred a longer “Rationale”. The latter is divided into two parts. The first provides a physical basis to the teleological character of quantum algorithms and, more in general, the building up of quantum correlation (in quantum algorithms, between the setting and the solution of the problem). The second discusses the possibility that the physical basis of the teleological character of quantum algorithms also applies to the evolutions of the living.
@@ 1.1 The physical basis of the teleological notion in the quantum domain
We must begin with an example of quantum speedup. The context, of course, is problem-solving. We consider the following very simple problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem; if it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the optimal quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case.
@@
First, it is logically equivalent to the problem-solver always knowing in advance, before she begins her problem-solving action, that the ball is in one of two given drawers and benefiting of this knowledge to solve the problem by opening just one drawer. Indeed, if the ball is in the drawer opened, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. Of course, what Alice should know in advance is the mutually exclusive $or$ of the three possible ways of pairing the number of the drawer with the ball with that of another drawer. Note that this also means knowing in advance half of the information that specifies the solution of the problem she will produce and measure in the future.
@@
The above immediately generalizes to any number, $N$, of drawers. Grover algorithm yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case222 The four drawer instance of Grover algorithm gives the solution with certainty; with more than four drawers, by “Grover algorithm” we will always mean Long’s version of it $\left[2\right]$, which always yields the solution with certainty. . Also now, this quantum speedup can be logically justified by the fact that Alice always knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is (in a mutually exclusive way) in one of the possible tuples of $\sqrt{N}$ drawers. She can then solve the problem in the best possible logical way by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
@@
Let us now move from logic to physics. Of course, the current quantum description of quantum algorithms does not describe any causal loop. Furthermore, being the customary “causal” quantum description of evolutions, where causality always goes in the same time-direction, it could not. Consistently with the present line of thought, we must conclude that this quantum description is incomplete.
@@
(i) First, it is a claim with an illustrious precedent, although apparently forgotten by the scientific community. We mean the 1935 argument of Einstein et al on the incompleteness of the quantum description $\left[3\right]$. We will see that the incompleteness of the quantum description of quantum algorithms is exactly the same lamented by Einstein et al for quantum nonlocality; moreover, the way to complete the quantum description of quantum algorithms also applies to quantum nonlocality and provides the hidden variables needed in both cases.
@@
We require that the determination of the problem-setting and thus that of the corresponding solution, instead of being entirely ascribed to the initial measurement of the problem-setting, is evenly shared between the initial measurement of the problem-setting and the final measurement of the corresponding solution. Since there is a plurality of mutually exclusive ways of evenly sharing, at the end we should take their quantum superposition.
@@
(i) each time symmetrization instance consists of a time-forward propagation (by a unitary transformation) going from the outcome of the initial measurement to the outcome of the final measurement and a time-backward propagation (by the inverse of the previous unitary transformation) going from the outcome of the latter measurement until replacing the previous outcome of the initial measurement. One can see that this is a causal loop that changes its own past.
@@
Also note that the time-symmetrization we are dealing with is the only possible way of completing the ordinary quantum description of the quantum algorithm in a way that describes the causal loops logically implied by the sheer existence of its quantum speedup. In fact, to describe a causal loop (any one) we must properly share the selection of the problem-setting between initial and final measurements, with the part of the selection performed by the initial measurement propagating forward in time and that performed by the final measurement backward in time. To obtain exactly the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, this sharing must be even. This exactly identifies the time-symmetrization we have performed. It also shows that this time-symmetrization is mandatory: it is the only way of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup and that the quantum algorithm must host.
Further below we will show that all the above holds for a large class of quantum algorithms, those solving oracle problems. Note that, so completed, the quantum description of quantum algorithms says much more than their ordinary quantum description; indeed, it provides the following unified quantitative description of the quantum speedups:
@@
Let us call the above statement in italics the advanced knowledge rule. This rule had already been achieved in the retrocausal approach to the quantum speedup of ourselves and coauthors $\left[4-8\right]$. In $\left[7\right]$, we had shown that it quantitatively explains the quantum speedup of all the major quantum algorithms. However, in these works, the time-symmetrization of the ordinary quantum description of the quantum algorithm was justified only by the fact that it turns out to produce, indeed, a unified quantitative description of the quantum speedups.
Now, this time-symmetrization has been shown to be necessary. It is the necessary consequence of the fact – we think inescapable – that the sheer existence of a quantum speedup over what is classically possible logically implies the mutual exclusive $or$ of well defined causal loops that must be implicit in the quantum algorithm. This should be a significant advancement. Causal loops, hitherto foreign to quantum mechanics, would instead be common in it, at least as common as quantum algorithms are.
In Section 2.2 we will see that the above advanced knowledge rule generalizes in the following rule:
@@
In the first place, the rule in question says that the building up of quantum correlation between two initially unrelated observables can always occur with quantum speedup. Of course, the notion of advanced knowledge (on the part of the problem-solver) of half of the information about the solution would lose any meaning outside problem-solving. However, note that it can be replaced by that of a corresponding reduction of the dimension of the Hilbert space in which the search for one-to-one correlation between the two observables takes place. This search can therefore occur with quantum speedup (see Section 2.2 for further detail).
In the second place, the causal loop rule of quantum correlation also applies to the case that the unitary evolution between the two one-to-one correlated measurement outcomes does not build up any quantum correlation between the two observables because they are maximally entangled to start with. This is the case of quantum nonlocality, where there is one-to-one quantum correlation between the outcomes of measuring (in the same basis) two maximally entangled observables in a state of spatial separation between the two respective subsystems.
Here, the unitary evolution between the two one-to-one correlated measurement outcomes must go, first, backward in time from the outcome of the first measurement to the time when the two subsystems were not yet spatially separate and then (in causal order) forward in time to the outcome of the second measurement. Of course, the eigenvalue selected by either measurement is not changed by the unitary process of spatial separation between the two subsystems.
As shown in $\left[8\right]$, the causal loops generated by time-symmetrizing the ordinary quantum description of the unitary evolution in question also locally go from the initial to the final measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell the not yet measured observable the state to assume in view of its future measurement. Therefore, these causal loops exactly play the role of the hidden variables that Einstein and the others $\left[3\right]$ envisaged to locally explain quantum nonlocality. Note that, in Einstein’s deterministic view of the selection of a measurement outcome, whether the two one-to-one correlated measurement outcomes are determined by the initial or the final measurement, or by the two in an even way, is indifferent. In other words, the hidden variables explanation of quantum nonlocality suggested Einstein and the others is compatible with the present notion of the causal loops, the latter being the hidden variables those authors were looking for.
@@
By the way, we do not know why the diagnosis of the incompleteness of the quantum description made by Einstein has been removed by the scientific community despite Einstein’s authority. Perhaps, because a satisfactory completion of it has never been provided. Of course, the price to pay for the present way of completing it is opening to retrocausality, a feature still avoided by mainstream quantum mechanics.
With this, we end the first part of the work.
@@
As we have seen, completing the ordinary quantum description of quantum algorithms also highlights their teleological character, that is, their being evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (the solution again). We can say that, in the quantum world, we have found the missing physical basis for the notion of teleological evolution.
This, naturally, raises the question whether the physical basis in question also applies to the evolutions of the living to which the teleological notion was once applied. The second part of the work addresses this question.
Incidentally, being thought without physical basis, and specially because of its finalistic character, the teleological view is usually classified among the beliefs. This, however, should not conceal the fact that, to start with, it was motivated by obvious empirical reasons. Indeed, it fits the functional structures and the evolutions of the living much better than the causal view. Of course, besides being why it was conceived in the first place, this is why, in relatively recent times, whether to abandon it has long been debated among biologists (see subsection 3.4.).
@@
This very enormity suggests another possibility. It is to adopt quantum cosmology $\left[9\right]$, namely the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way. We could then think of scaling the quantum features down from the largest and longest-lasting evolution in existence, that of the universe.
In fact, a consequent quantum interpretation of the Anthropic Principle $\left[10-13\right]$ would provide a solution to our problem. It would say that the evolution of the universe toward the “value” of life (in particular, “existence of sentient life” or “absence of life”) builds up a quantum correlation between the value (for values) of the fundamental physical constants since the beginning of time and the value of life today; this evolution could consequently occur with quantum speedup and would thus be teleological in character.
@@
Let us see this in more detail. As is well known, the Fine-tuned Universe version of the Anthropic principle is a way of expressing the fact that the slightest change of the value of the fundamental constants since the beginning of time would have produced a radically different universe most probably unable to accommodate life.
@@
This version of the principle just asserts that there is a correlation (we assume for simplicity one-to-one) between the value of the fundamental constants since the beginning of time and the value of life today. Of course, this correlation must have been built up by the evolution of the universe under the fundamental laws and constants. In fact, at the beginning of time, the value of the fundamental constants and that of life were unrelated: no matter the value of the fundamental constants, there could have been no life.
@@
Under the assumption that the evolution of the universe is unitary quantum, the correlation asserted by the anthropic principle becomes quantum, and we have seen that the building up of quantum correlation can always occur with quantum speedup. Indeed, the unitary building up of quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today is formally identical to a quantum search algorithm that unitarily builds up a quantum correlation between the value of the problem-setting since its beginning and the value of the solution at its end – one can see that this is a way of viewing what quantum algorithms do.
@@
The evolution of the universe toward the value of life can therefore occur with quantum speedup and consequently be teleological in character – the “can occur” is said in the sense that there is certainly the room for its occurring with a quantum speedup444 Note that, in problem-solving, advanced knowledge on the part of the problem-solver of half of the information about the solution does not imply a quantum speedup but means room for it; in order that there is a quantum speedup, the quantum algorithm must benefit of this advanced knowledge. .
@@
In the universe we are in, what can occur with quantum speedup is of course the evolution of the universe toward life. This room for quantum speedup is then inherited by the evolution of life on Earth, even if decohered. In fact, first, evolution times must be the same for the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, and its image in the evolution of the universe. Second, the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
@@
In the end, being able to occur with quantum speedup, the evolution of life on Earth can be teleological; the same would consequently hold for the evolutions of life on Earth, in particular those of species. Since it is reasonable, and conservative, to assume that the search for the value of life by the evolution of the universe under the fundamental physical laws and constants is random, its room for quantum speed should become an actual quantum speedup, as in quantum random search algorithms $\left[14\right]$.
@@
Section 2. is about the teleological notion in the quantum framework; it can be seen as an extension and formalization of the first part of the Rationale.
Section 3. is about the present, scientific, teleological notion in natural evolutions – it is an extension of the second part of the Rationale.
Section 4 provides the prospects for further research; Section 5. tries a positioning of the work; Section 6. yields the conclusions.

### v13

@@ Abstract
1 Introduction
This work proposes a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the very goal it will produce in the future. It is, of course, a puzzling notion since it implies a causal loop. For over two millennia, it has been applied by natural philosophy first to the functional structure of the living (Aristotle: teeth were created for the purpose of chewing) and then, with the advent of the notion of evolution of the living, to the latter. It was then dismissed with the advent of modern science because of its alleged lack of physical basis. Of course, this is because the teleological notion implies retrocausality, even today a very controversial if not ostracized notion – certainly a notion never accepted by mainstram science.
What we think should change this (centuries-old) position of science is the advent of quantum mechanics and, more specifically, the relatively recent discovery of the quantum computational speedup. In fact, we will show that quantum algorithms are scientific examples of teleological evolutions: it is their teleological character that yields their quantum computational speedup over the corresponding classical algorithms.
Also with coauthors, we have developed the retrocausal explanation of the quantum speedup $\left[1-5\right]$. In it, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information about the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps (oracle queries) than the minimum required in the classical case. The number of computation steps needed in the optimal quantum case is that needed to solve the problem in an optimal logical way benefiting of the advanced knowledge in question. By the way, more precisely, what the problem-solver knows in advance is the mutually exclusive $or$ of all the possible halves of the information about the solution.
In the previous works, we showed that that explanation fits all the major quantum algorithms, see $\left[4\right]$. In the present work, we provide an elementary derivation of it, truly elementary and perfectly checkable, one just had to think about it. We analize Grover’s quantum search algorithm $\left[6\right]$ from the standpoint of classical logic (we then generalize to oracle quantum computing). This is legitimate: involving only the measurements of commuting observables (the initial measurement of the setting of the problem and the final measurement of the corresponding solution), quantum algorithms are submitted to classical logic. In a few lines, we show that the sheer existence of the quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of the above mentioned causal loops.
Of course, this retrocausal explanation of the quantum speedup is a unifying quantitative one. Given an oracle problem (any one), it provides the number of computation steps (oracles queries) needed to solve it in an optimal quantum way; it is the number needed to solve it in an optimal logical way benefiting of the advanced knowledge of one of the possible halves of the information about the solution. Let us call this the advanced knowledge rule.
Note that, despite its obvious convenience, this rule is absent from mainstream quantum information, which relies on the notion of causal evolution, i.e. an evolution where causality always goes forward in time. Here, the only way of knowing the number of computation steps (oracle queries) needed to solve an oracle problem in an optimal quantum way is finding the quantum algorithms that does it. Of course, this is an object of high ingenuity, very uncertain achievement, and different from oracle problem to oracle problem. As far as we know, excluding extremely simple quantum algorithms, this object has been reached in only one case, that of Grover’s quantum search algorithm.
There is, however, a catch. To date, no work involving retrocausality, although published in mainstream physics journals, sometimes by famous authors, has ever made it into mainstream science.
There is also the opposite “however”. However, dealing with the quantum speedup, the retrocausality we are dealing with presents a substantial novelty with respect to all the forms proposed before. Indeed, it has two very significant pro’s:
(a) One is that it is logically implied by the sheer existence of the quantum speedup in an elementary and perfectly checkable way (as we will soon see).
(b) The other is that it turns out to be just an interpretation of the unitary interplay between quantum superpositions and interferences that characterizes the quantum computation of the solution of a problem. In other words, suitable unitary interplays between quantum superpositions and quantum interferences – the two pillars of quantum mechanics – generate retrocausal effects. Retrocausality is implicit in mainstream quantum mechanics!
Note that, despite being just an interpretation of the current quantum description of quantum algorithms, the retrocausality we are dealing with is a retrocausality in the fullest sense of the world. Knowing in advance half of the information about a future measurement outcome means information sent backward in time from the future to the past, full fledged retrocausality indeed.
We would argue that the two above pro’s should guarantee to retrocausality, at last, right of citizenship in mainstream quantum mechanics. We believe that this is the main result of the present work. Let us have have, anyway, a summary of the rest. We will see that everything fits together.
We now move from the logic of quantum algorithms to their physics.
Of course, their current quantum description, a causal one (i.e. with causality always going forward in time), does not and cannot describe causal loops. We must conclude that it is incomplete, of course in a fundamental way. However, first, we should get rid of a trivial incompleteness also present in the current quantum description of quantum algorithms. Being limited to the unitary computation of the solution (where the quantum speedup resides) and the final measurement of the latter, their current quantum description is trivially incomplete for the lack of the initial measurement. This is also trivially emended, by adding to their current quantum description that of the upstream process of setting the problem. An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random. We call this extension of the current quantum description of the quantum algorithm its ordinary quantum description, “ordinary” because it is still the causal quantum description of mainstream science.
We will show that the only possible way of completing it, by making it to describe the causal loops logically implied by the sheer existence of the quantum speedup, is by time-symmetrizing it. This leaves the unitary part of the ordinary quantum description of the quantum algorithm mathematically unaltered by changes the behavior of causality along it. The single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
This completion of the ordinary quantum description of quantum algorithms immediately generalizes to that of quantum correlation (in quantum algorithms, between the setting and the solution of the problem). More specifically, it extends to any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue of the observable selected by the initial measurement (e. g., the problem-setting is not changed by the unitary computation of the solution).
By the way, this is also the case of quantum nonlocality, that is, the measurement (in the same basis) of two maximally entangled observables in a state of spatial separation between the two respective subsystems. Time-symmetrizing its ordinary quantum description, exactly as we did with that of quantum algorithms, yields a quantum superposition of causal loops that play the role of the hidden variables envisaged by Einstein and the others in their celebrated EPR paper $\left[7\right]$. They locally tell the not yet measured observable the state to assume in view of its future measurement.
All the above also says that quantum algorithms are full fledged teleological evolutions, indeed, evolutions toward a goal, the solution of the problem, with an attractor in the very goal they will produce in the future, the solution of the problem again.
Of course, showing that quantum algorithms are scientifically justified teleological evolutions raises the question whether this should change the position of science on natural teleological evolutions.
It should be noted that whether to abandon the teleological view of the functional structures and the evolutions of the living has long been debated among evolutionary biologists and is sometimes discussed also today. This is because it is a view that fits the living things in a very natural way. Therefore, we believe that checking whether the present scientific notion of teleological evolution is applicable also to the natural sciences should be a real must.
What we found would seem to be significant. It is that a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle, we would say “very surprisingly”, yields that the evolution of the universe (under the fundamental physical laws and constants) toward the “value of life” (either “existence of sentient life” of “absence of life”) is formally identical to a quantum search algorithm. In either case, the evolution could occur with quantum speedup and thus be teleological in character; in either case, it should occur with quantum speedup even if the search is random. Of course, the “quantum cosmological assumption” $\left[8\right]$ is indeed an assumption, that is, is completely hypothetical in character. However, the long-standing claim that the teleological notion is devoid of any possible physical basis would in any case be falsified: indeed, a possible physical basis does exist.
@@ 2 Rationale
It should already be clear from the introduction that this work is unconventional. In effect, it is a different way of approaching quantum information. Therefore, before going to the formalizations and further detail, in the present “Rationale” we provide a conceptual presentation of the work.
The Rationale is divided into two parts. The first provides a physical basis to the teleological character of quantum algorithms and, more in general, the building up of quantum correlation (in quantum algorithms, between the setting and the solution of the problem). The second examines the possibility that the physical basis of the teleological character of quantum algorithms also applies to the evolutions of the living.
@@ 2.1 The physical basis of the teleological notion in the quantum domain
We must begin with an example of quantum speedup. The context, of course, is problem-solving. We consider the following very simple problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem; if it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the optimal quantum case – by the optimal quantum search algorithm of Grover $\left[6\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case.
@@
Here, by “opening only one drawer”, we mean performing only one oracle query. Of course, only one drawer is opened but for a quantum superposition of all the possible drawer numbers. In the following, we analyze this behavior of the quantum algorithm from the standpoint of logic, classical logic of course. Indeed, involving only the measurements of commuting observables (the problem-setting and the corresponding solution of the problem) quantum algorithms are submitted to classical logic (there is no special “quantum logic” involved here).
First, the fact of always solving the problem by opening just one drawer is logically equivalent to the problem-solver always knowing in advance, before she begins her problem-solving action, that the ball is in one of two given drawers and benefiting of this knowledge to solve the problem by opening just one drawer. Indeed, if the ball is in the drawer opened, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. Of course, what Alice should know in advance is the mutually exclusive $or$ of the three possible ways of pairing the number of the drawer with the ball with that of another drawer. Note that this also means knowing in advance half of the information that specifies the solution of the problem she will produce and measure in the future.
@@
The above immediately generalizes to any number, $N$, of drawers. Grover algorithm yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case222 The four drawer instance of Grover algorithm gives the solution with certainty; with more than four drawers, by “Grover algorithm” we will always mean Long’s version of it $\left[9\right]$, which always yields the solution with certainty. . Also now, this quantum speedup can be logically justified by the fact that Alice always knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is (in a mutually exclusive way) in one of the possible tuples of $\sqrt{N}$ drawers. She can then solve the problem in the best possible logical way by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
@@
Let us now move from logic to physics, of course quantum physics (the quantum algorithm indeed). Naturally, the current quantum description of quantum algorithms does not describe any causal loop. Furthermore, being the customary “causal” quantum description of evolutions, where causality always goes in the same time-direction, it could not. Consistently with the present line of thought, we must conclude that this quantum description is incomplete.
@@
(i) First, it is a claim with an illustrious precedent, although apparently forgotten by the scientific community. We mean the 1935 argument of Einstein et al on the incompleteness of the quantum description $6$. We will see that the incompleteness of the quantum description of quantum algorithms is exactly the same lamented by Einstein et al for quantum nonlocality; correspondingly, the way to complete the quantum description we will provide for quantum algorithms, will also work for quantum nonlocality providing the hidden variables – here the causal loops – needed in both cases.
@@
We require that the determination of the problem-setting and thus that of the corresponding solution, instead of being entirely ascribed to the initial measurement of the problem-setting, is evenly shared between the initial measurement of the problem-setting and the final measurement of the corresponding solution. Since there is a plurality of (logically) mutually exclusive ways of evenly sharing, at the end we should take their quantum superposition.
@@
(i) each time symmetrization instance consists of a time-forward propagation (by a unitary transformation) going from the outcome of the initial measurement to the outcome of the final measurement and a time-backward propagation (by the inverse of the previous unitary transformation) going from the outcome of the latter measurement until replacing the previous outcome of the initial measurement. One can see that this is a non-trivial causal loop, namely one that changes its own past.
@@
Also note that the time-symmetrization we are dealing with is the only possible way of completing the ordinary quantum description of the quantum algorithm in a way that describes the causal loops logically implied by the sheer existence of its quantum speedup. In fact, with the kind of process we are dealing with, to describe a causal loop (any one), we must properly share the selection of the problem-setting between initial and final measurements, with the part of the selection performed by the initial measurement propagating forward in time and that performed by the final measurement backward in time. To obtain exactly the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, this sharing must be even. This exactly identifies the time-symmetrization we have performed. It also shows that this time-symmetrization is mandatory: it is the only way of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup and that the quantum algorithm must host.
Further below we will show that all the above holds for a large class of quantum algorithms, those solving oracle problems. Note that, so completed, the quantum description of quantum algorithms says much more than their ordinary quantum description; in fact, it highlights the following unifying quantitative feature of the quantum speedups:
@@
Let us call the above statement in italics the advanced knowledge rule. This rule had already been achieved in the retrocausal approach to the quantum speedup of ourselves and coauthors $\left[1-5\right]$. In $\left[4\right]$, we had shown that it quantitatively explains the quantum speedup of all the major quantum algorithms. However, in these works, the time-symmetrization of the ordinary quantum description of the quantum algorithm was justified only by the fact that it turns out to produce, indeed, a unified quantitative description of the quantum speedups.
Now, this time-symmetrization has been shown to be necessary. It is the necessary consequence of the fact – we think inescapable – that the sheer existence of a quantum speedup over what is classically possible logically implies the mutual exclusive $or$ of well defined causal loops that must be implicit in the quantum algorithm. This should be a significant advancement. Causal loops, hitherto foreign to mainstream quantum mechanics, would instead be common in it, at least as common as quantum algorithms are.
We open a parenthesis to explain why we provided a review of a retrocausal approach to the quantum speedup that has already been published. The point is that the approach in question is probably little known as it never entered mainstream physics. The latter is a fate common to most publications dealing with retrocausality, starting with the famous1949 one by Wheeler and Feynman $\left[10\right]$. Although, at the time, the latter gave rise to several discussions, none of it ever entered mainstream physics. Of course, the idea that the effect can precede the cause can be difficult to accept. However, we have already seen in the Introduction, that this effect preceding the cause must be implicit in quantum mechanics as it can always be replaced by a suitable, unitary, interplay between quantum superpositions and interferences, of course the two the pillars of quantum mechanics.
Back to our previous line of thought, in Subsection 3.2, we will see that the need of completing the quantum description by time-symmetrizing it is not limited to quantum algorithms. It extends as follows:
@@
In the first place, this rule says that the unitary building up of one-to-one quantum correlation between two initially unrelated observables that does not change the eigenvalue of the observable selected by the initial measurement can always occur with quantum speedup (see Subsection 3.2 for further detail). Note that, if the process is not a quantum algorithm, the notion of advanced knowledge on the part of the problem-solver of half of the information about the solution must be replaced by that of a corresponding reduction of the dimension of the Hilbert space in which the search for one-to-one correlation between the two observables takes place.
In the second place, the rule also applies to the case that the unitary evolution does not build up any quantum correlation between the two observables because they are maximally entangled to start with. This is the case of quantum nonlocality, where there is one-to-one quantum correlation between the outcomes of measuring (in the same basis) two maximally entangled observables in a state of spatial separation between the two respective subsystems. Of course, the eigenvalue selected by either measurement is not changed by the unitary process of spatial separation between the two subsystems.
As shown in $\left[5\right]$, the causal loops generated by time-symmetrizing the ordinary quantum description of quantum nonlocality locally go from the initial to the final measurement outcome via the time the two subsystems were not yet space-separate, where they locally tell the not yet measured observable the state to assume in view of its future measurement. Therefore, these causal loops exactly play the role of the hidden variables that Einstein and the others $\left[7\right]$ envisaged to locally explain quantum nonlocality.
@@
By the way, we do not know why the diagnosis of the incompleteness of the quantum description made by Einstein and the others has been removed by the scientific community despite Einstein’s authority. Perhaps, because a satisfactory completion of it has never been provided. Of course, the price to pay for the present way of completing it is opening to retrocausality, a feature still avoided by mainstream quantum mechanics.
With this, we end a first part of the Rationale.
@@
As we have seen, completing the ordinary quantum description of quantum algorithms highlights their teleological character, that is, their being evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (the solution again). We can say that, in the quantum world, we have found the missing physical basis for the notion of teleological evolution.
This, naturally, raises the question whether the physical basis in question also applies to the evolutions of the living to which the teleological notion, before the scientific revolution, was applied. The second part of the Rationale addresses this question.
Incidentally, being thought without physical basis, and specially because of its finalistic character, the teleological view has been classified among the beliefs. This, however, should not conceal the fact that, to start with, it was motivated by obvious empirical reasons. Indeed, the teleological view fits the functional structures and the evolutions of the living much better than the causal view. Of course, besides being why it was conceived in the first place, this is why, in relatively recent times, whether to abandon it has long been debated among biologists (see Subsection 4.4.).
@@
However, the very enormity of this distance suggests a possible solution. It is to adopt quantum cosmology $\left[8\right]$, namely the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way. We could then think of scaling the quantum features down from the largest and longest-lasting evolution in existence, that of the universe.
Let us show how a consequent quantum interpretation of the Anthropic Principle $\left[11-14\right]$, the Fine-tuned Universe version of it, could provide a solution to our problem.
@@
As is well known, this version of the Anthropic Principle is a way of expressing the noticeable fact that the slightest change of the value of the fundamental physical constants since the beginning of time would have produced a radically different universe most probably unable to accommodate life.
@@
This version of the principle just asserts that there is a correlation (we assume for simplicity one-to-one) between the value (for values) of the fundamental constants since the beginning of time and the “value of life” today (in particular, “existence of sentient life” or “absence of life”). Of course, this correlation must have been built up by the evolution of the universe under the fundamental laws and constants. In fact, at the beginning of time, the value of the fundamental constants and that of life were unrelated: no matter the value of the fundamental constants, there could have been no life.
@@
Under the assumption that the evolution of the universe is unitary quantum, the correlation asserted by the anthropic principle becomes quantum, and we have seen that this kind of building up of quantum correlation (unitary and never changing then eigenvalue selected by the initial measurement) can always occur with quantum speedup (see Subsection 3.2 for further detail). Indeed, the unitary building up of quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today is formally identical to a quantum search algorithm that unitarily builds up a quantum correlation between the value of the problem-setting since its beginning and the value of the solution at its end – one can see that this is a way of viewing what quantum algorithms do.
@@
The evolution of the universe toward the value of life can therefore occur with quantum speedup and consequently be teleological in character – the “can occur” is said in the sense that there is certainly the room for its occurring with a quantum speedup444 Note that, in problem-solving, advanced knowledge on the part of the problem-solver of half of the information about the solution does not imply a quantum speedup but means room for it; of course, in order that there is a quantum speedup, the quantum algorithm must benefit of this advanced knowledge. .
@@
In the universe we are in, what can occur with quantum speedup is the evolution of the universe toward life. This room for quantum speedup is then inherited by the evolution of life on Earth, even if decohered. In fact, first, evolution times must be the same for the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, and its image in the evolution of the universe. Second, the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
@@
In the end, being able to occur with quantum speedup, the evolution of life on Earth can be teleological; the same would consequently hold for the evolutions of life, in particular those of species. Since it is reasonable, and conservative, to assume that the search for the value of life by the evolution of the universe under the fundamental physical laws and constants is random, its room for quantum speed should become an actual quantum speedup, as in quantum random search algorithms $\left[15\right]$.
@@
Section 3. is about the teleological notion in the quantum framework; it is the formalization of the first part of the Rationale.
Section 4. is about the present, scientific, teleological notion in natural evolutions – it is an extension and formalization of the second part of the Rationale.
Section 5. provides the prospects for further research; Section 6. tries a positioning of the work; Section 7. yields the conclusions.